# Space as Time Through Neuron Position Learning

### v1

@@ Abstract
Biological neural networks exist in physical space where distance determines communication delays: a fundamental space-time coupling absent in most artificial neural networks. While recent work has separately explored spatial embeddings and learnable synaptic delays in spiking neural networks, we unify these approaches through a novel neuron position learning algorithm where delays relate to the Euclidean distances between neurons. We derive gradients with respect to neuron positions and demonstrate that this biologically-motivated constraint acts as an inductive bias: networks trained on temporal classification tasks spontaneously self-organize into local, small-world topologies with modular structure emerging under distance-dependent connection costs. Remarkably, we observe unprompted functional specialization aligned with spatial clustering without explictly enforcing it. These findings lay the groundwork for networks in which space and time are intrinsically coupled, offering new avenues for mechanistic interpretability, biologically inspired modelling, and efficient implementations.
@@ 1 Introduction
Brain networks face fundamental trade-offs between metabolic costs and information processing capabilities. Networks must minimise the costs of building and maintaining connections in physical space while optimising for computational capability. This constraint shapes brain organisation across species and may explain why brains converge on similar structural solutions (Achterberg et al., 2023), including sparse, small-world (Van den Heuvel et al., 2016) architectures and modular (Kaiser & Hilgetag, 2004) organisation.
@@
To analyse the role of space, recent studies have started focusing on the spatial embedding of biological neural networks (Achterberg et al., 2023; Erb et al., 2025; Sheeran et al., 2024; Vasilache et al., 2025) – a critical dimension largely overlooked in computational models. However, these studies do not investigate the relationship between this spatial embedding and time – despite the spatial locations of neurons and their relative timing of information processing being naturally intertwined in the brain. Biological neural networks exist in physical space, where distance translates directly into time through conduction delays—a ‘space as time’ relationship (Voges & Perrinet, 2010; Izhikevich & Hoppensteadt, 2009) that is fundamental to brain organisation. The cost of connections is proportional to their length in three-dimensional space, and conduction delays vary dramatically across the nervous system due to differences in path length, diameter, and can even be optimised through myelin plasticity (Bonetto et al., 2021), i.e the brain could implement something akin to delay learning.
@@
One of the first algorithms that employed a form of delay learning used multapses with various delays between neuron pairs and a corresponding trainable weight (Bohte et al., 2002). Since the delays could not be explicitly optimised, this algorithm could be understood more as a form of structure learning where delays are fixed and the optimal values are found through optimising the connectivity weights. The relationship between network structure and delays is still of interest, as Hammouamri et al. highlighted the benefits of delay learning in sparse networks, and Mészáros et al. (2024) observed that structural plasticity in networks with delays leads to better performance than explicit delay learning. Regardless, recent methods focus on gradient-based approaches (Hammouamri et al., ; Göltz et al., 2024; Sun et al., 2023), with Mészáros et al. (2024) introducing delay learning for recurrent connections.
@@
The intertwinement of space and time creates natural constraints on both the structure and the dynamics of neural networks. Spatial clustering minimises costly long-range connections while creating fast communication within local regions and slower communication between distant areas. Thus, spatial organisation inherently generates temporal separation, where different brain regions operate on different timescales determined by their connectivity patterns, yielding hierarchically organised areas with increased temporal receptive windows (Chang et al., 2022). Evolution has exploited this relationship, as exemplified by the barn owl’s sound localisation system, which relies on precisely tuned delays arising from spatial circuit organisation (Carr & Konishi, 1988; Ghosh et al., 2025). These observations have not yet been modelled as, in most cases, neural networks discard spatial and temporal information. Even Spiking Neural Networks (SNNs) typically only adapt weights and biases – failing to exploit the spatial embedding and delay structure that biological systems have optimized over evolution.
Achterberg et al. (2023) proposed an approach for studying spatial structure in ANNs – giving us a way to connect them to temporal processing. However, their models only saw the cost of the connections, rather than the benefit of their structure. Erb et al. (2025) and Vasilache et al. (2025) explored position learning through distance-based synaptic weights, but disregarded the temporal aspects of space.
@@
Now the foundations for both spatial aspects (through spatially embedding neurons) and the temporal aspects (through delay learning) have been laid, a natural next step is combining these into a single framework. Here, we do just that and derive a gradient-based learning algorithm for SNNs, where synaptic delays are determined by the Euclidean distance between two connected neurons. Furthermore, we derive the gradients of loss functions with respect to the positions of neurons, introducing a neural network shape learning algorithm. We study the various effects of the algorithm through training recurrent SNNs on the Spiking Heidelberg Digits (SHD) (Cramer et al., 2020) classification task. We find that position learning networks rely more on local computations and form circuits associated with particular ‘tasks’ after training. Not only do our findings have implications for our understanding of the brain but also, suggest new ways of studying neural networks and potential directions for future neuromorphic implementations.
@@ 2.1 Position learning links structure to task
We first introduce a novel neuron position learning algorithm, in which neurons move in space so as to minimise the objective function. This makes use of the fact that delays between neurons can augment performance. Intuitively, a neuron with a great spatial separation should therefore correspond to a longer delay. Recently, various delay learning algorithms have been derived, and shown to be an effective way of improving the performance of SNNs (Hammouamri et al., ; Sun et al., 2023; Göltz et al., 2024; Mészáros et al., 2025). Mészáros et al. (2025) was the first to introduce delay learning for recurrent connections. Spatial embeddings become particularly interesting with bidirectional connectivities, thus we derived our equations in a recurrent framework. In a 2-dimensional space, we define the delay between neuron $j$ and $i$ as:
@@
where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ denote the positions of neurons $i$ and $j$ in a 2-dimensional space. Thus, the spike arrival time at neuron $m$ becomes:
$t_{lm}\equiv t_{l}+d_{mn(l)}=\\ t_{l}+\sqrt{(x_{m}-x_{n(l)})^{2}+(y_{m}-y_{n(l)})^{2}},$ (2)
where $n(l)$ is the neuron, that fired the $l$-th spike, at time $t_{l}$. Similarly to the original algorithm (Wunderlich & Pehle, 2021) and its extensions (Mészáros et al., 2025), we also work with Leaky Integrate-and-Fire (LIF) neurons, where each neuron has an input current $I$, and a membrane voltage $V$.
@@
From (Mészáros et al., 2025), equation (6), we know
$\frac{d\mathcal{L}}{dd_{ji}}=-w_{ji}\sum_{t_{k}\in\,\parbox{18.49411pt}{\scriptsize spikes from\ $i$}}(\lambda_{I}-\lambda_{V})_{j}\Big|_{t_{k}+d_{ji}}.$ (4)
and analaguosly for $\frac{d\mathcal{L}}{dd_{ij}}$ and by taking the derivative of equation (1), we get
@@
Taken together,
@@
Here, $\mathcal{L}$ is the loss and $w_{ij}$ is the synaptic weight. $\lambda_{I}$ and $\lambda_{V}$ denote the backwards sensitivities corresponding to the input current and membrane voltage, respectively. We observe that these updates are a sum of of all presynaptic and postsynaptic delay gradients, normalised by $(x_{i}-x_{j})/d_{ji}$. We thus arrive at a learning algorithm with the same computational benefits as before – neurons have the same computational requirements for the forward and backward passes, the costly synaptic updates only happen at spike times in both directions.
Temporally complex tasks benefit from introducing delays (Hammouamri et al., ; Habashy et al., 2024; Sun et al., 2023; Göltz et al., 2024) and adding delays to recurrent connections is particularly beneficial for small networks (Mészáros et al., 2025; Queant et al., 2025). This raises the question: do spatial embeddings serve as a useful inductive bias by defining delays as the Euclidean distance between two neurons? If so, what would otherwise be considered a constraint could be leveraged.
We investigate this by training an SNN with learnable positions on the SHD classification task (Cramer et al., 2020). By learning positions instead of synaptic delays, we reduce the number of free parameters from $n^{2}$ to $n\times d$, where $n$ is the number of neurons, and $d$ is the number of dimensions. Furthermore, this set-up yields delay distributions which inherit the following properties from the Euclidean distance:
  • Zero diagonals: Autapses take effect in the next timestep
  • Symmetry: Delay from neuron $i$ to neuron $j$ is the same as from $j$ to $i$
@@
To ascertain that position learning is useful, we first compare position learning networks of various sizes and embeddings of various dimensions, against networks with non-spatially constrained synaptic, and axonal delays, and networks without any delays. We could have additionally compared against networks with fixed random delays, but it raises the question of how to initialise these delays. In recurrent connections, this is a particularly complex and understudied question. We illustrate the comparisons of the networks in the Appendix.
With a fixed number of trainable parameters, we observe that networks with delays always outperform networks without them. As network size (i.e. number of hidden neurons) increases, the performance gaps decrease. This is not surprising, as all architectures have recurrent connections, which, similar to delays, can serve as a temporal memory. As the network gets more and more neurons to store information in, delays become less and less necessary. Spatially embedded networks perform similarly to networks with unconstrained delay learning. Axonal delays, however, perform at least as well as the synaptic delays derived from the Euclidean distances. Note, that the error bars are large because we study small networks, to make sure that there is a clear benefit of delays.
The formulation does not allow for neurons to be placed on top of each other, but in all cases we initialise the shape such that distances are close to 0 (and for non-spatial architectures we initialise all delays to 0).
We conclude that recurrent delay learning for positions improves task performance – a necessity for the rest of our studies. From here, we focus on one network size and embedding dimension – 2D networks with a hidden layer size of 128 – since we observed no fundamental differences between network sizes and embedding dimensions.
Similarly to Achterberg et al. (2023) we also study the effects of L1 regularisation, and L1 regularisation scaled with neuron distance. For our studies, we force neurons into an n-dimensional sphere (so that the effect of the length-scaled regularisation will be constrained), with a diameter of 62, such that the maximum possible delay is equal to the maximum delay allowed on Loihi 2 (Davies et al., 2021). Since regularisation will not drive weights to become exactly 0, we tune our hyperparameters based on validation accuracies obtained after dropping the 90% lowest magnitude connections.
@@ 2.2 Nearby neurons become functionally connected
We looked at the relationship between neuron positions in the hidden layer and the inputs they preferentially respond to, post-training. First, we tried to predict the X, Y positions of hidden neurons from their corresponding 700-dimensional input to hidden weight vectors using ridge regression. We found that – based on the observed R2 scores shown in Table 1 – we can do this quite accurately, suggesting that neurons with similar input patterns are physically close as well.
Init No reg. L1 L1+dist.
X $0.07$ $0.82\pm 0.1$ $0.72\pm 0.1$ $0.6\pm 0.1$
Y $0.07$ $0.78\pm 0.1$ $0.59\pm 0.1$ $0.61\pm 0.1$
Table 1: R2 scores
We also looked at the centre of mass for each input neuron, by taking the mean of all neuron positions weighted with the corresponding synaptic strength, and, as figure 1 shows, neurons in both networks clearly separate depending on which ‘part’ of the input they focus on (i.e higher or lower frequencies).
Figure 1: A Input example from the SHD dataset. B Example of preferred positions of input neurons based on the learnt synaptic strengths. C Example of preferred position of input neurons, based on their activity. D Examlpe of preferred position over time. The depicted plots are from networks trained with L1+distance cost regularisation, but we observe such spatial patterns under all training circumstances.
Next, we measured the average (over input samples) spiking activity of each hidden neuron in a 5 ms window after each input spike and then took the average of this 3-dimensional tensor (input neuron $\times$ hidden neuron $\times$ window) over time. Note that this measure is different from measuring firing rates, as postsynaptic spikes need to happen in a small window after presynaptic spikes. Doing this, we again end up with a matrix that is the same shape as the input-to-hidden weight matrix. We can now again take the weighted average of neuron position according to these values. For results, see Figure 1. Instead of taking the average over time, we can also measure the average sensitivity for each timestep and look at where computation is happening for each input neuron. We observe that over time, computation ”moves” towards the centre and spreads out again by the end of the trial.
2.3 Higher modularity and lower entropy through distance cost
Next we studied various metrics in trained networks. Since the weights are regularised, we can expect many weights to be close to but not exactly zero. Thus, we measure the weighted modularity metric (for definition see Appendix A). In Figure 2, we can observe that the peak modularity Q is achieved relatively early on in training, and the distance cost needs to be introduced, for higher modularity (i.e position learning will not inherently lead to more modular weight matrices).
Shannon entropy (for definition see Appendix A) quantifies the amount of unpredictability. A highly predictable system has low entropy and will exhibit a degree of clustering around particular weight values. An unpredictable system has high entropy and will be more uniform in its distribution. While the previously introduced modularity metric Q shows the extent to which the network can be partitioned into subcommunities, it does not measure the degree of uncertainty of the weight distribution. In Figure 2 we observe, that position learning networks achieve lower entropy than non-spatial learning networks, but introducing the distance cost decreases entropy even further.
In this work, we are interested in the effect of space on network structure, and thus only introduced the distance cost to the regularisation term. Achterberg et al. (2023) and Sheeran et al. (2024) also introduced a communicability cost, which we did not employ here. However, we can still measure the communicability entropy, similarly to Sheeran et al. (2024) (for definition see Appendix A). In figure 2, we can observe that networks with the introduced distance cost yield a lower entropy. Given that weight and communicability entropy are both lower for positions learning networks, we can conclude that these architectures yield a small number of highly communicable connections, with an increasingly ordered topology.
We also measured the small-worldness on the connectivity matrix with the 90% lowest magnitude connections dropped. Again, we observe that the distance cost term increases the metric. Interestingly, without the distance cost, we observe lower small-worldness in position learning networks, compared to synaptic delay learning.
We can also visualise the learnt network shapes together with the corresponding synaptic strengths and figure 2 compares these with and without the spatial cost. We can observe that the network shapes end up looking similar (with neurons moving close to the perimeter), but without the spatial cost, we end up having stronger long-range connections. There is no direct driving force on the neuron positions, and thus they try to maximise the possible delays, but through the distance-based regularisation, the connectivity patterns are more local.
Figure 2: Network shapes without (A) and with spatial cost (B). Line thickness corresponds to synaptic strength. Neurons move close to the perimeter in both cases, but when introducing the distance cost, we have fewer long strong connections. C-F Training dynamics over epochs. D shows that the network with the introduced cost term achieves a higher modularity than the other two. In E we observe that spatial networks achieve a lower Shannon entropy in their weights, regardless of the cost term. On C we see that if calculate the Shannon entropy on the communicability matrix derived from the weight matrix, the network with the distance cost has a lower value. On C we observe that introducing the cost term yields higher small-worldness.
In this section, we have applied the metrics used in Sheeran et al. (2024) to our models. We observe the same distinctions between the non-spatial and spatial architectures, but to a smaller extent than what they showed. While it is possible that this is due to the introduction of position learning and delays, it is more likely that it is thanks to the fact that we tuned our hyperparameters for high accuracy (after pruning), and weaker regularisation. Furthermore, comparing networks with and without a distance cost with the same regularisation strength is not a fair setup. As neurons can change positions, this could too heavily influence the network dynamics, as the regularisation term is continuously changing not only according to the weights, but according to the network shape as well. Additionally, Sheeran et al. (2024) and Achterberg et al. (2023) added a communicability matrix to the regularisation term, which we do not study here.

### v2

@@ Abstract
Biological neural networks exist in physical space where distance influences communication delays: a fundamental coupling between space and time absent in most artificial neural networks. While recent work has separately explored spatial embeddings and learnable synaptic delays in spiking neural networks, we unify these approaches through a novel neuron position learning algorithm where delays relate to the Euclidean distances between neurons. We derive gradients with respect to neuron positions and demonstrate that this biologically-motivated constraint acts as an inductive bias: networks trained on temporal classification tasks spontaneously self-organize into local, clustered topologies and a modular, efficiently wired structure emerges if connection costs are distance-dependent. Remarkably, we observe functional specialization aligned with spatial clustering without explicitly enforcing it. These findings lay the groundwork for networks in which space and time are intrinsically coupled, offering new avenues for mechanistic interpretability, biologically inspired modelling, and efficient implementations.
Spiking Neural Networks, Delay Learning, Spatial Embeddings
@@ 1 Introduction
Figure 1: (A) A light cone represents the trajectory through spacetime of a flash of light emitted from a single point at a single instant, expanding outward in all directions over time. Similarly, in our framework, neurons’ spikes (depicted as stars) “propagate” through space as they evolve in time, forming conical regions of influence in spacetime. Where these cones intersect (highlighted with a red line), the accumulated input can become strong enough to drive a neuron at that spatial location (c) to spike. (B) We train a single hidden layer Recurrent Spiking Neural Network on Spiking Heidelberg Digits for classification. (C) Compared to the standard approach, we not only train the synaptic weights, but the neuron positions as well, by defining synaptic delays as Euclidean distances. (D) Illustration of how individual synaptic delay gradients influence position learning updates. Neurons sum all pre- and post-synaptic delay gradients and normalise them according to the distance from the other neuron. (E) Training neuron positions in this framework shows the benefit of long-distance connections, compared to previous approaches, which only focused on the cost of connections. We observe that networks grow during training, with some neurons diverging more and others staying closer to each other. The lines depict the trajectory of neuron positions throughout training, and the points the final positions neurons converged on. The colours of trajectories illustrate the time course of the neuron positions throughout training.
@@
Brain networks face fundamental trade-offs between metabolic costs and information processing capabilities. Networks must minimise the costs of building and maintaining connections in physical space while optimising for computational capability. This constraint shapes brain organisation across species and may explain why brains converge on similar structural solutions (Achterberg et al., 2023), including sparse, clustered architectures (Van den Heuvel et al., 2016) and modular organisation (Kaiser and Hilgetag, 2004).
@@
Biological neural networks exist in physical space, where distance translates directly into time through conduction delays – a ‘space as time’ relationship (Voges and Perrinet, 2010; Izhikevich and Hoppensteadt, 2009; Keller et al., 2024) that is fundamental to brain organisation. Recent studies have started considering the spatial embedding of biological neural networks (Achterberg et al., 2023; Erb et al., 2026; Sheeran et al., 2024; Vasilache et al., 2025; Tohouri, 2025), focusing on the cost of long connections, rather than also considering delays and their potential computational benefits. These studies do not investigate the relationship between the spatial embedding and time – even though the relative location of neurons and transmission delays between them are strongly related in the brain.
@@
Delay learning in Spiking Neural Networks (SNNs) has recently been shown to increase the capacity of networks, i.e. to allow for compressing networks in terms of their size and their parameter precision (Sun et al., 2025). One of the first algorithms that employed a form of delay learning used multapses with various delays between neuron pairs and a corresponding trainable weight (Bohte et al., 2002). Since the delays were not explicitly optimised, this algorithm can be understood more as a form of structure learning where delays are fixed and the optimal values are found by optimising the connectivity weights. The relationship between network structure and delays remains an active research area. Hammouamri et al. (2024) highlighted the benefits of delay learning in sparse networks, and Mészáros et al. (2024) observed that structural plasticity in networks with delays leads to better performance than explicit delay learning. Regardless, other recent works focus on gradient-based approaches (Hammouamri et al., 2024; Göltz et al., 2025; Sun et al., 2023), with Mészáros et al. (2025) introducing delay learning for recurrent connections.
@@
The intertwinement of space and time creates natural constraints on both the structure and the dynamics of neural networks. Spatial clustering minimises costly long-range connections while creating fast communication within local regions and slower communication to distant areas. Evolution has exploited this relationship, as exemplified by the barn owl’s sound localisation system, which relies on precisely tuned delays arising from spatial circuit organisation (Carr and Konishi, 1988; Ghosh et al., 2025).
@@
Here, we present the natural next step: combining spatial embedding and delay learning into a single framework. Instead of learning synaptic delays directly, we define delays as the Euclidean distance between the connected neurons and derive the gradient of our loss function with respect to neuron positions. Gradient descent with respect to neuron positions effectively introduces a neural network shape learning algorithm. Within this framework, long connections, which would simply represent a high cost in previous works (Achterberg et al., 2023; Erb et al., 2026; Sheeran et al., 2024; Vasilache et al., 2025; Tohouri, 2025), can now also be harnessed for computation. We study the various effects of the algorithm using recurrent SNNs trained on the Spiking Heidelberg Digits (SHD) (Cramer et al., 2020) classification task. We find that networks with position learning rely predominantly on local computations and form circuits associated with particular ‘tasks’ after training. Not only do our findings have implications for our understanding of the brain, but they also suggest new ways of studying neural networks and potential directions for future neuromorphic implementations.
@@ 2.1 Position learning links structure to task
Figure 2: (A) $R^{2}$ score for ridge regression of hidden neurons’ positions from their input connectivity weights. Regression has high $R^{2}$ for all networks, and particularly for spatial networks with strong distance-dependent cost regularisation. We depict the mean of the results over 20 random seeds, error bars represent standard deviation.(B) We depict the ”preferred target positions” of input neurons by weighting hidden neurons’ positions with the corresponding input-to-hidden weights (second row of C and D), or the correlations between input and hidden spikes (third row of C and D). (C) Effects of training (from left to right) in a network trained with weak L1 regularisation. First row: learnt network shapes; neurons spread out with training. Second row: preferred input target positions derived from weights. We see a clear separation of preferred target positions for the different frequency bands in the input. Last row: preferred target positions derived from activity (correlations). The positions appear to form a trajectory. (D) Same as C for a network with strong regularisation. Measuring the preferred target positions using the weights now also looks more like trajectories instead of clusters.
Figure 3: Preferred positions over time in an example network. Positions move towards the centre and spread out to the edges through the trial.
Figure 4: (A) Network shapes with and without distance cost and increasing regularisation strength from left to right. Line thickness corresponds to synaptic strength. Networks increase in size with stronger regularisation, until they reach a breaking point, and become small again. (B) Delay distributions weighted with their corresponding synaptic strength over 20 random seeds. The increase of network sizes can also be observed through the distributions.
We first introduce a novel neuron position learning algorithm, in which neurons move in space to minimise the objective function, which depends on neurons’ position through synaptic delays that are defined by the Euclidean distance of neurons. Mészáros et al. (2025) were the first to introduce delay learning for recurrent connections, and, since spatial embeddings become particularly interesting with bidirectional connectivities, we derived our equations in this framework. In a two-dimensional space, we define the delay between neuron $i$ and $j$ as:
@@
where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ denote the positions of neurons $i$ and $j$. Thus, the spike arrival time for spike $t$ at neuron $j$ becomes:
$t_{j}\equiv t+d_{ji}=t+\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}},$ (2)
where $i$ is the neuron that fired at time $t$, which arrived at neuron $j$ at time $t_{j}$. Similar to the original algorithm (Wunderlich and Pehle, 2021) and its extensions (Mészáros et al., 2025), we also work with Leaky Integrate-and-Fire (LIF) neurons, where each neuron has an input current $I$, and a membrane voltage $V$.
@@
From Mészáros et al. (2025), equation (6), we know
$\frac{d\mathcal{L}}{dd_{ji}}=-w_{ji}\sum_{t_{k}\in\,\parbox{18.49411pt}{\scriptsize spikes from\ $i$}}(\lambda_{I}-\lambda_{V})_{j}\Big|_{t_{k}+d_{ji}}$ (4)
where $w_{ij}$ is the synaptic weight; and $\lambda_{I}$ and $\lambda_{V}$ denote the backwards sensitivities corresponding to the input current and membrane voltage, respectively. Next, by taking the derivative of equation (1), we get
@@
Taken together with the expression for $\frac{d\mathcal{L}}{dd_{ij}}$, obtained by rewriting (4), we obtain:
@@
We investigate our algorithm by training SNNs with learnable positions on the SHD classification task (Cramer et al., 2020). By learning positions instead of synaptic delays, we reduce the number of free parameters from $n^{2}$ to $n\times d$, where $n$ is the number of neurons, and $d$ is the dimensionality of the space neurons are embedded in. Furthermore, this set-up yields delay distributions which inherit the following properties from the Euclidean distance:
  • Zero diagonals: Spikes arriving through self connections are immediate
  • Symmetry: The delay from neuron $i$ to neuron $j$ is the same as from $j$ to $i$
@@
To ascertain that position learning is useful, we first compared the classification accuracy of position learning networks and standard networks with and without delays. Recurrent delay learning through positions performs similarly to non-spatial networks with fewer trainable parameters (supplementary figure 7 in the Appendix). For the remainder of this paper, we focus on analysing 2D networks with a hidden layer size of 128 since we observed no fundamental differences between network sizes and embedding dimensions.
In the following, we study unconstrained synaptic delay learning (Non-spatial) and position learning (Spatial) networks with L1 regularisation on the hidden weights, as well as a spatially weighted L1 regularisation (Spatial+cost) in which each synaptic weight is penalised proportionally to the distance between neurons (Achterberg et al., 2023). In the following, we study unconstrained synaptic delay learning (Non-spatial) and position learning (Spatial) networks with L1 regularisation on the hidden weights, as well as a spatially weighted L1 regularisation (Spatial+cost), in which each synaptic weight is penalised proportionally to the distance between neurons (Achterberg et al., 2023). In the latter case, we use the normal L1 penalty and apply its gradient, scaled by distance, to the hidden weight updates. Note that this implies that L1 regularisation is only applied to weights, and not to positions. This distance-dependent penalty can be interpreted as a resource constraint that discourages both excessive connectivity and long-range connections, reflecting limitations on wiring and metabolic cost. Since the distances between neurons change dynamically during training, when calculating the spatially weighted L1 regularisation, we normalise the distances by dividing by the mean. This ensures that the overall strength of the regularisation term remains comparable to standard L1 regularisation and does not dominate the loss function solely due to changes in network geometry.
@@ 2.2 Nearby neurons become functionally connected
First, we focus on the relationship between the computational role of individual neurons and their spatial position. We looked at the relationship between neuron positions in the hidden layer and the inputs they preferentially respond to after training. First, we predicted the $x$, $y$ positions of hidden neurons from their corresponding 700-dimensional input to hidden weight vectors using ridge regression. We found that – based on the observed $R^{2}$ scores shown in Figure 2A – that we can do this accurately, suggesting that neurons’ input patterns are strongly correlated to their position. We observe an increase in $R^{2}$ scores as regularisation strengthens.
To get an understanding of where computations are happening, we also looked at the ‘centre of mass’ of each input neuron’s projections, by taking the mean of all hidden neuron positions weighted by the corresponding input to hidden synaptic strength (Figure 2B, equation 7 in the Appendix). As the examples in Figure 2C show, after optimising neuron positions, input neurons representing similar audio frequencies appear to project to specific locations. We also observe that, as regularisation strength increases, preferred target locations for input neurons representing similar frequencies not only become more localised, but they even organise in to structures that look like trajectories (Figure 2D).
Next, we measured the influence of input neurons on hidden layer activity by analyzing evoked hidden layer spikes. We measured the average spiking activity (across input samples) of each hidden neuron in the 5 timesteps following each input spike (for definition see equation 8 in the Appendix). We then took the average of this 3-dimensional tensor (input neuron $\times$ hidden neuron $\times$ 5) over the 5 timesteps. This process again results in a matrix with the same shape as the input-to-hidden weight matrix. Therefore, we can again calculate the ‘centre of mass’ of each hidden neuron’s influence based on these values. Figure 2C and D show examples of this process, and depicting preferred target positions in this manner shows the formation of clear trajectories, regardless of regularisation strength.
Importantly, both clustered and trajectory-like patterns indicate non-random spatial organisation. While clustering reflects the localisation of inputs with similar frequencies to nearby hidden neurons, the trajectory-like structures show a continuous spatial organisation of input frequency in the hidden layer.
Figure 5: Regularisation effects on topology. We depict the mean of the results over 20 random seeds, error bars represent standard deviation. (A) Weighted modularity. ‘Spatial+cost’ network shows a monotonic increase, while the modularity of other networks seems to peak with weaker regularisation. (B) Weighted clustering coefficient. Resource constraints increases clustering in all cases, but plain L1 regularisation has a stronger effect. Spatial networks without the distance cost achieve particularly high clustering. (C) Weighted wiring efficiency. Following our intuition, efficiency is particularly high with a distance cost.
Instead of taking the average over time of the previously measured ‘sensitivities’, we can also measure the average sensitivity at each timestep and observe where computation is happening for each input neuron. As shown in Figure 3, early in the trial, preferred positions are concentrated towards the spatial centre of the network where neurons are closer to each other and therefore primarily connected via short synaptic delays. Later on, preferred positions spread outward towards more peripheral regions, the interactions between which rely on longer delays due to the increased distances. This temporal progression suggests a shift from fast, short-delay interactions supporting rapid feature extraction early in the trial to longer-delay interactions later on, enabling information to be maintained and integrated over extended timescales. Such behaviour is consistent with the notion of temporal hierarchies in neural computation (Moro et al., 2024).
2.3 Relative importance of short and long delays under constraints
As expected, increasing L1 weight regularisation strength results in lower accuracy and higher sparsity (see Figure 8 in the appendix). Figure 4A visualises the shapes of example networks with increasing regularisation strengths both with and without the distance cost. As regularisation strength increases, both the network with and without spatial cost increase in size (with the long distance synapses being stronger in the absence of the distance cost). However, at extreme regularisation strength, we reach a breaking point, where networks become very small again. Note that, for this regularisation strength, the non-spatial network and the spatial network without the distance cost converge on solutions that do not rely on the recurrent connections, i.e. all recurrent connections can be pruned without accuracy loss. We speculate that as the networks have fewer weights (with increased weight regularisation), they have to depend on longer and longer delays. However, while long delays can be introduced on individual connections in the non-spatial architecture, in spatial networks, this can only be achieved by moving neurons further apart, which increases all delays associated with these neurons. Thus, extreme resource constraints show the importance of short delays. Observing the synaptic strength-weighted delay distributions (figure 4B) reveals that across architectures and regularisation strengths, delays seem to be concentrated in the short range.
2.4 Learnt topology is constraint-dependent
Next, we studied the topology of the learnt network under increasingly strong regularisation. In Figure 5A, we can observe that, as the regularisation strength increases, the network with the additional distance cost achieves higher and higher modularity (for definition see equation (12) in the Appendix). In the other networks, while modularity increases to some extent, it reaches its maximum observed value with weaker regularisation than the maximum studied strength.
We also calculated a weighted clustering metric from our connectivity matrices, using a shuffled weight matrix as the null-model (for definition see equation (13) in the Appendix). Interestingly, we see a rapid increase for the non-spatial and spatial network, particularly compared to the network with distance cost. However, again, stronger regularisation does not increase the value further. In networks with the distance cost we observe a similar trend, but with lower values.
Although most connections in biological neural networks are short – reflecting spatial and energetic constraints – such networks typically contain a small but important set of long-distance connections. Thus, the wiring cost is higher than a network rewired for minimum cost, but much lower than a network rewired for maximum cost (Kaiser and Hilgetag, 2006). Thus, in our networks, we measure how the wiring cost relates to the minimum and maximum possible wiring cost (see equation (20) in the Appendix for definition). We observe that our spatial networks are more efficiently wired than non-spatial architectures and show higher efficiency at higher regularisation strengths. Intuitively, strong distance-scaled regularisation yields particularly efficient wiring.